If an employee deposits Rs. 2,000 at the end of each year into his company’s plan which pays 7% interest compounded quarterly, how much will he have in the account at the end of 5 years?
1
Expert's answer
2019-05-22T08:55:16-0400

In this case result be next:

If we have simple percents (interest compounded for base deposit), we will receive:

S=S0(1+npl/(100m))+nR(1+(n-1)pl/(200m))={S0=2000, m=4, n=5,l=4,p=7}=2000*(1+5*7*4/(100*4))+5*2000*(1+4*7*4/(200*4))=14100 Rs.

If we have complex percents (interest compounded for each new value of the deposit), we will receive:

S=S0*(1+p/(100*m))n*l+R(1-(1+p/(100*m))nl)/(1-(1+p/(100*m))l)={S0=2000, m=4, n=5,l=4,p=7}=

=2000*(1+7/(100*4))5*4+2000 (1-(1+7/(100*4))5*4)/(1-(1+7/(100*4))4)=14373.78